OpenStudy (ledah):

Solve by factoring: x^2+5x-6=0

OpenStudy (anonymous):

first, think of 2 numbers that multiply together to make 6, and add together to make 5.

OpenStudy (ledah):

2 and 3

OpenStudy (anonymous):

no search more its 6 and 1 now you relate them

OpenStudy (ledah):

.... what? 2 times 3=6 and 2+3 =5. those are both right what does 6 and 1 have to do with it?

OpenStudy (anonymous):

6 time of 1 is also 6 but for a specific combination of 6 and 1 the sum will be 5 and product will be -6 try now

OpenStudy (ledah):

Confusing me....

OpenStudy (anonymous):

(x+6)*(x-1)

OpenStudy (anonymous):

well okay. 2 and 3. now you have to think up of combinations of those numbers to give positive 5, and negative 6. But 2 and 3 don't have such a combination. So you need to think of another 2 numbers that can give that. so what @apoorwa said, with 1 and 6. now think, for positive 5, and negative 6, should the 1 be negative, or the 6.

OpenStudy (anonymous):

x+6=0 x= -6 x-1=0 x = 1

OpenStudy (anonymous):

x^2+6x-x-6=0 x(x+6)-1(x+6)=0 (x+6)(x-1)=0

OpenStudy (anonymous):

an alternate way, which apoorwa did above, is to split the middle - so the 5x into two. then it's all a matter of taking the common factor out. the thing in the bracket should be the same. so that'll constitute as one of the factors of x, and the thing outside the bracket, you can just collect, and that'll be the other factor. :)

OpenStudy (ledah):

meaning what?